OpenStudy (anonymous):

One last question: What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

I know I have to use the elimination method but how?

OpenStudy (anonymous):

multiply (3x + 2y = 8) by co-efficient of x or y multiply (2x + 3y = 2) by co-efficient of x or y you get 6x + 4y = +16 + 6x + 9y = +6 - - - (sign changed for subtraction) ------------------------- -5y = 10 y = -2 put y in any equation to get x = 4

OpenStudy (anonymous):

Are you 100% positive? If so thank you. : )

OpenStudy (anonymous):

multiply equation 1 by 2 and eq 2 by 3 (coeff of x).

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

pleasure.